# Golden Touchstone: A Comprehensive Bilingual Benchmark for Evaluating Financial Large Language Models

### v1

@@ 2.1 Financial Large Language Models
In recent years, large language models (LLMs) tailored for the financial domain have become increasingly important. Several significant FinLLMs have been developed: BloombertGPT (Wu et al., 2023) marks the beginning of the FinLLM era. FinGPT (Yang et al., 2023) is an open-source framework designed for financial LLMs. It focuses on a data-centric approach, offering transparent resources for developing FinLLMs. The framework highlights the importance of an automated data curation pipeline and lightweight low-rank adaptation techniques. The follow-up work, FinGPT-Forecaster, continues to use LLM for stock price prediction, which is not ideal for numerical computation tasks. PIXIU (Xie et al., 2023) provides a comprehensive framework that includes the first financial LLM fine-tuned on LLaMA with instruction data, a 136K instruction dataset for fine-tuning, and an evaluation benchmark with 5 tasks and 9 datasets. This framework supports the evaluation and development of financial LLMs. CFGPT (Li et al., 2023) introduces a Chinese Financial Generative Pre-trained Transformer framework. It consists of CFData, a dataset for pre-training and supervised fine-tuning; CFLLM, a financial LLM capable of processing financial texts; and CFAPP, a deployment framework for real-world financial applications. DISC-FinLLM Chen et al. (2023) is based on a Multiple Experts Fine-tuning Framework. It enhances general LLMs by adding multi-turn question answering, domain text processing, mathematical computation, and retrieval-enhanced generation capabilities. XuanYuan 2.0 (Zhang and Yang, 2023) is the largest Chinese chat model to date, built on the BLOOM-176B architecture. It introduces a novel training method called hybrid-tuning to address catastrophic forgetting, combining general-domain and domain-specific knowledge. FinMem (Yu et al., 2023), FinRobot (Yang et al., 2024c), and FinReport (Li et al., 2024) predict stock price trends using agents and tools, such as quantitative models that use multiple alpha factors and alternative factor models that analyze news and textual data. FinVis-GPT (Wang et al., 2023b) and FinTral Bhatia et al. (2024) develop multi-modal financial models, similar to other visual language models (VLM) (Zhu et al., 2023; Liu et al., 2023; Lu et al., 2023b). These models represent the cutting-edge in FinLLMs, each contributing unique approaches and advancements to the field of AI in finance.
@@ 2.2 Benchmarks for FinLLMs
In the current landscape of open-source financial evaluation benchmarks, FLUE (Shah et al., 2022) was the first open-source benchmark introduced for financial language understanding in English. It includes five financial tasks: Financial Sentiment Analysis (Malo et al., 2014; Maia et al., 2018b), Financial News Classification (Sinha and Khandait, 2021), Structure Boundary Detection (FinSBD3, 2021), Named Entity Recognition (Alvarado et al., 2015), and Question Answering (Maia et al., 2018a). Building on FLUE, FinGPT (Yang et al., 2023) introduced the Financial Relation Extraction task with the addition of the FinRED dataset (Sharma et al., 2022), while also expanding the sentiment analysis task by incorporating TFNS (Magic, 2022) and NWGI (Yang, 2023) datasets. Additionally, FinGPT created prompt-based instructions across all datasets, culminating in the Financial Instruction Tuning Benchmark. Around the same time, other open-source benchmarks, such as PIXIU (Xie et al., 2023) and FinBen (Xie et al., 2024), were developed. PIXIU includes a wide range of financial tasks and datasets, while FinBen, expanding on PIXIU, introduced numerous supplementary financial datasets to evaluate LLMs comprehensively and gauge their proficiency across various financial scenarios.In the Chinese domain, BBT-Benchmark (Lu et al., 2023a) was the first Chinese financial benchmark, covering multiple financial NLP tasks, including sentiment analysis, news classification, summarization, relation extraction, and question answering. Later, FinEval (Zhang et al., 2023) benchmark emerged, providing an evaluation framework for LLMs in financial contexts, encompassing multiple-choice questions across finance, economics, accounting, and certification. DISC-FinLLM (Chen et al., 2023) evaluated financial LLMs on top of these benchmarks, introducing additional tasks focused on financial calculations and current events to assess LLMs in applied financial scenarios. The most recent Chinese Financial benchmark, CFBenchmark (Lei et al., 2023), introduced three key tasks like Financial Entity Recognition, expanding evaluation resources for Chinese financial tasks.
@@
3 Benchmark Design
3.1 Current Benchmark Status
In current open-source financial benchmarks, FLUE(Shah et al., 2022) is among the most widely used benchmarks for English financial NLP tasks, encompassing five financial tasks: financial sentiment analysis, news headline classification, named entity recognition, structural boundary detection, and question answering. FinGPT(Yang et al., 2023) extended this benchmark with additional financial tasks. PIXIU(Xie et al., 2023) proposed a more comprehensive benchmark, covering more financial tasks and datasets than FLUE and introducing a stock movement prediction task. Building on PIXIU, FinBen(Xie et al., 2024) added numerous financial datasets to thoroughly evaluate large language models (LLMs) and assess their proficiency across diverse financial scenarios. In the Chinese financial domain, BBT-Benchmark(Lu et al., 2023a) was the first Chinese benchmark for financial evaluation, covering financial news classification, summarization, relation extraction, question answering, negative news detection, and sentiment classification for financial social media texts. Additionally, the FinEval(Zhang et al., 2023) benchmark provides an evaluation framework for LLM knowledge in finance, including multiple-choice datasets across finance, economy, accounting, and certification fields. Subsequently, the FinanceIQ(Zhang and Yang, 2023) and CFBenchmark(Lei et al., 2023) benchmarks were introduced, with FinanceIQ including a broader range of Chinese financial multiple-choice datasets, while CFBenchmark focuses on three major financial task categories: financial entity recognition, financial text classification, and financial content generation.
However, data quality issues in certain tasks within these benchmarks—characterized by inconsistent, unstructured input information—often hinder the accurate comprehension and response generation of financial LLMs. Large language models currently face challenges in numerical understanding and computation, which may be attributed to multiple factors such as tokenizers, positional encoding, and others (Shen et al., 2023; Akhtar et al., 2023; Schwartz et al., 2024). Therefore, using specialized agents is often a more ideal approach for handling these tasks rather than relying on vanilla large language models. This issue is especially evident in PIXIU’s credit analysis, question answering, and stock movement prediction tasks, where model performance metrics are frequently low. Additionally, varying label formats across datasets for the same task make it challenging to match model outputs with labels directly. Notably, most current financial benchmarks are monolingual, lacking bilingual evaluation options to assess financial models comprehensively in both English and Chinese.
Table 1: Diversity of Financial Analysis Tasks Across Different Financial Large Language Model Benchmarks
FinGPT-Bench (Wang et al., 2023a) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
FinBen (Xie et al., 2024) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
BBT-Fin (Lu et al., 2023a) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Fin-Eval (Zhang et al., 2023) $\checkmark$
FinanceIQ (Zhang and Yang, 2023) $\checkmark$
CFBenchmark (Lei et al., 2023) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Table 2: Language Coverage, Systematicity, Adaptability, and Model Training Stage for Benchmarks. Systematicity refers to whether benchmarks are established according to a comprehensive system standard. Adaptability indicates whether the tasks are suitable for large language models.
Benchmarks Language Systematicity Adaptability Model Training
EN CN Cont. Pre-train Instr. Tuning
FinGPT-Bench (Wang et al., 2023a) $\checkmark$ Medium High $\checkmark$
FinBen (Xie et al., 2024) $\checkmark$ High Medium $\checkmark$
BBT-Fin (Lu et al., 2023a) $\checkmark$ Medium High $\checkmark$
Fin-Eval (Zhang et al., 2023) $\checkmark$ High High
FinanceIQ (Zhang and Yang, 2023) $\checkmark$ Medium High $\checkmark$ $\checkmark$
CFBenchmark (Lei et al., 2023) $\checkmark$ High High $\checkmark$ $\checkmark$
Golden-Touchstone $\checkmark$ $\checkmark$ High High $\checkmark$ $\checkmark$
To address these limitations, we propose a unified benchmark that incorporates representative financial datasets in both English and Chinese. By selecting high-quality, authoritative datasets and optimizing inputs—such as using news-based input for stock trend prediction rather than tabular data—we aim to make tasks more compatible with LLM assessment. This bilingual benchmark will facilitate model evaluation on both English and Chinese financial tasks, align financial NLP tasks across languages, and provide a holistic measure of financial model performance across linguistic contexts.
The financial sub-tasks included in previous financial benchmarks are shown in Table.1, which illustrates the diversity of financial analysis tasks across different benchmarks. Table.2 provides an overview of the language coverage, systematicity, adaptability, and model training of these benchmarks. In this context, systematicity refers to whether benchmarks are established according to a comprehensive and standardized framework, while adaptability assesses whether the tasks are suitable for large language models. As seen in Table 1, the task coverage of earlier benchmarks is not fully comprehensive, with some financial tasks missing. To address this, we selected representative datasets from various open-source benchmarks based on the task categories in Table 1, ensuring a broad diversity of tasks in the Golden-Touchstone benchmark. Therefore, Golden-Touchstone includes tasks that broadly cover those in previous benchmarks. From Table 2, it is evident that most current financial benchmarks are monolingual, with inconsistent levels of systematicity and adaptability. In response, we have compiled the first bilingual financial benchmark from open-source datasets, achieving high standards in terms of both systematicity and adaptability.
3.2 Golden Touchstone Benchmark Design
Golden Touchstone benchmark redefines popular open-source financial benchmarks, as illustrated in Figure 2. We classify financial NLP tasks by type—financial NLU (Natural Language Understanding) and financial NLG (Natural Language Generation)—and by language, distinguishing between English and Chinese datasets. These dimensions guide the organization of high-quality open-source datasets. NLU tasks include Financial Sentiment Analysis, Financial Knowledge Examination, Financial Information Extraction, and Financial Text Understanding, while NLG tasks include Financial Summary Analysis and Financial Question Answering. Collectively, these cover eight sub-tasks: Sentiment Analysis, Classification, Entity Recognition, Relation Extraction, Multiple Choice, Summarization, Question Answering, and Stock Movement Prediction. For the English benchmark, tasks were selected primarily from the PIXIU(Xie et al., 2023) and FinBen(Xie et al., 2024) financial benchmarks, covering sentiment analysis, classification, entity recognition, question answering, stock movement prediction, summarization, and credit scoring. Due to low data quality and inconsistent labeling formats in some datasets, we excluded certain lower-quality data while retaining those with relatively higher quality. FinBen’s credit scoring task, a binary classification task, was merged into the general classification task. Additionally, as the stock movement prediction task in FinBen uses time-series tabular data, which tends to be ineffective for prediction in large language models (LLMs) due to LLMs’ limited sensitivity to numerical data in inference contexts, we replaced this task with the DJIA (Aaron7Sun, 2016) dataset, using daily news to predict Dow Jones Index trends. To ensure alignment with the Chinese benchmark tasks, we also incorporated the Financial Relation Extraction task from FinGPT (Yang et al., 2023) and the Multiple Choice task from the CPA-Eval (Yang et al., 2024b) dataset, using the validation set as the test set due to the absence of labels in the test set. For the Chinese benchmark, tasks were chosen primarily from the BBT-benchmark (Lu et al., 2023a) and FinEval benchmark (Zhang et al., 2023) , including sentiment analysis, classification, entity extraction, question answering, summarization, relation extraction, and multiple-choice tasks. As test sets for these benchmarks lack labels, we used the validation set for evaluation. To expand the multiple-choice tasks, we added the CPA (Yang et al., 2024b) dataset, using its validation set as a test set, also without labels. For alignment with the English benchmark tasks, we introduced stock movement prediction with the stockA (Zou et al., 2022) dataset, which includes individual stock news and stock data for five days post-publication. The price movement label is based on the difference between the final day’s closing price and the initial net asset value, enabling LLM-based predictions on stock trends informed by news. Dataset details are provided in Table.3 and Table.4.
@@ 3.2 Golden Touchstone Benchmark Design
Table 3: Overview of English Finance Evaluation Datasets by Task Type, Sample Sizes (Training, Validation, Test), and Evaluation Metrics
Task Dataset Train Valid Test Metrics
Sentiment Analysis FPB 3100 776 970 Weighted-F1 ACC
FiQA-SA 750 188 235 Weighted-F1 ACC
Classification Headlines 71900 10300 20500 Weighted-F1 ACC
FOMC 1984 - 496 Weighted-F1 ACC
lendingclub 9417 1345 2691 Weighted-F1 MCC
Entity Recognition NER 408 103 98 Entity-F1
Relation Extraction FinRE 27558 - 5112 Relation-F1
Multiple Choice CFA 1884 100 20 Weighted-F1 ACC
Summarization EDTSUM 8000 - 2000 ROUGE BLEU
Question Answering FinQa 6251 883 1147 RMACC
ConvfinQa 8890 2210 1490 RMACC
Stock Movement Prediction DJIA 1591 - 398 Weighted-F1 ACC

### v2

@@ 4.1 Financial Large Language Models
In recent years, large language models (LLMs) tailored for the financial domain have gained significant attention. BloombertGPT (wu2023bloomberggpt) marked the beginning of the FinLLM era. FinGPT (yang2023fingpt) introduced an open-source framework emphasizing a data-centric approach with lightweight low-rank adaptation techniques. PIXIU (xie2023pixiu) provided a comprehensive framework, presenting the first financial LLM fine-tuned on LLaMA with a 136K instruction dataset and evaluation benchmark. CFGPT (li2023cfgpt) developed a Chinese Financial Generative Pre-trained Transformer framework, encompassing dataset, model, and deployment capabilities. DISC-FinLLM (chen2023disc) enhanced general LLMs through a multiple experts fine-tuning framework, expanding domain-specific capabilities.
@@ 4.2 Benchmarks for FinLLMs
In the current landscape of open-source financial evaluation benchmarks, FLUE (shah2022flue) was the first open-source benchmark introduced for financial language understanding in English. It includes five financial tasks: Financial Sentiment Analysis (malo2014fpb; maia2018finqa_sa), Financial News Classification (sinha2021headlines), Structure Boundary Detection (finsbd3), Named Entity Recognition (alvarado2015NER), and Question Answering (FiQA_flue). Building on FLUE, FinGPT (yang2023fingpt) introduced the Financial Relation Extraction task with the addition of the FinRED dataset (sharma2022finred), while also expanding the sentiment analysis task by incorporating TFNS (tfns2022) and NWGI (nwgi2023) datasets. Additionally, FinGPT created prompt-based instructions across all datasets, culminating in the Financial Instruction Tuning Benchmark. Around the same time, other open-source benchmarks, such as PIXIU (xie2023pixiu) and FinBen (xie2024finben), were developed. PIXIU includes a wide range of financial tasks and datasets, while FinBen, expanding on PIXIU, introduced numerous supplementary financial datasets to evaluate LLMs comprehensively and gauge their proficiency across various financial scenarios.In the Chinese domain, BBT-Benchmark (lu2023bbt) was the first Chinese financial benchmark, covering multiple financial NLP tasks, including sentiment analysis, news classification, summarization, relation extraction, and question answering. Later, FinEval (zhang2023fineval) and FinanceIQ (zhang2023xuanyuan)benchmark emerged, providing an evaluation framework for LLMs in financial contexts, encompassing multiple-choice questions across finance, economics, accounting, and certification. DISC-FinLLM (chen2023disc) evaluated financial LLMs on top of these benchmarks, introducing additional tasks focused on financial calculations and current events to assess LLMs in applied financial scenarios. The most recent Chinese Financial benchmark, CFBenchmark (lei2023cfbenchmark), introduced three key tasks like Financial Entity Recognition, expanding evaluation resources for Chinese financial tasks.
@@
5 Conclusion
6 Limitations
Despite these advancements, the benchmark currently exhibits certain limitations, including a limited range of NLG tasks and a focus solely on single-modality. Future enhancements will include the integration of additional NLG tasks, such as extended text generation for financial report analysis and more sophisticated sentiment assessments. Dataset biases and model scalability are also challenges. Furthermore, we plan to expand the benchmark to cover other financial sectors such as insurance and futures trading, thus broadening the scope and applicability of financial LLM assessments across diverse scenarios. Also, the performance of Touchstone-GPT on specific tasks within the Golden Touchstone benchmark, particularly in stock market prediction, requires further improvement. Our subsequent research will explore the incorporation of agent-based and retrieval-augmented generation (RAG) methods to augment the model’s capabilities in numerical computation and real-time news analysis. Additionally, we aim to venture into multimodal modeling, integrating visual data and time-series data for tasks such as financial time-series forecasting, financial chart analysis, and content generation.
7 Ethical Statements
We do not see our work to have possible harmful outcomes. We follow the ACL ethical guidelines when conducting the research in this paper.
@@ 7 Ethical Statements
Appendix A Overview of Finance Evaluation Datasets by Task Type, Sample Sizes (Training, Validation, Test), and Evaluation Metrics
In the Table.4 and 5, we show the overview of finance evaluation datasets by task type, sample sizes (training, validation, test), evaluation metrics and license.
Table 4: Overview of English Finance Evaluation Datasets by Task Type, Sample Sizes (Training, Validation, Test), Evaluation Metrics and License
Task Dataset Train Valid Test Metrics License
Sentiment Analysis FPB 3100 776 970
Weighted-F1
ACC
CC BY-SA 3.0
FiQA-SA 750 188 235
Weighted-F1
ACC
MIT License
Classification Headlines 71900 10300 20500
Weighted-F1
ACC
CC BY-SA 3.0
FOMC 1984 - 496
Weighted-F1
ACC
CC BY-SA 3.0
lendingclub 9417 1345 2691
Weighted-F1
MCC
Entity Recognition NER 408 103 98
Entity-F1
CC BY-SA 3.0
Relation Extraction FinRE 27558 - 5112
Relation-F1
MIT License
Multiple Choice CFA 1884 100 20
Weighted-F1
ACC
CC BY-SA 3.0
Summarization EDTSUM 8000 - 2000
ROUGE
BLEU
MIT License
Question Answering FinQa 6251 883 1147
RMACC
MIT License
ConvfinQa 8890 2210 1490
RMACC
MIT License
Stock Movement Prediction CMIN-US 88297 9010 8480
Weighted-F1
ACC
MIT License